The sum of two numbers is 38. The larger number is 6 more than the smaller number. What are the numbers?
Answer by macston(5194)

x= the smaller number
x+(x+6)=38
2x+6=38
2x=32
x=16 ANSWER 1: The smaller number is 16.
x+6=16+6=22 ANSWER 2: The larger number is 22.
CHECK:
Sum of two numbers is 38
16+22=38
38=38